I am stuck can you help me. Thank-you
1. Write a quadrtic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0
Solution:10. only solution
The equation is _=0.
2.Write a quadratic equation in the variable x having given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0
-sqrt7.8sqrt7
The equation is_=0.

Answer by stanbon(57984)

1. Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0
Solution:10. only solution
If 10 is the only solution it must have multiplicity two.
Since 10 is a solution, (x-10) is a factor.
The equation is (x-10)^2 = 0
x^2 - 20x + 100 = 0
------------------------------------------------------
2.Write a quadratic equation in the variable x having given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0
-sqrt7.8sqrt7
(x+sqrt(7))(x-8sqrt(7)) = 0
The equation is x^2 - 7sqrt(7)x - 56 = 0
================================================
Cheers,
Stan H.